A normal breath takes in about 1.0 L of air. Assum
Answer #1

Q4.

V = 1L of air

MW of air = 28.8 g/mol

D = 0.97 g /L

find molecules in each breath

1 mol = 6.022*10^23 molecules

calculate mass per breath:

Density = mass/Volume

mass = D*V = 0.97 g/L * 1 L = 0.97 grams of air per breath

1 mol of air = 28.8 g

therefore

0.97 g --> 0.97/28.8 = 0.0336805 mol of air

and we know that

1 mol of air = 6.022*10^23 molecules

so

0.0336805 mol of air = (0.0336805)(6.022*10^23) = 2.028239*10^22 molecules of air per breath

nearest answer is A)

2.0*10^22